if the sides of a square are increased by 4 inches, the area becomes 100 square inches. Find the length of the sides of the original square.
Answer by jim_thompson5910(35256)

(x+4)^2 = 100

x+4 = sqrt(100)

x+4 = 10

x = 10-4

x = 6

So the length of the sides of the original square is 6 inches.